OpenStudy (anonymous):

what is the solution to the equation log4 x - log4 2=3?

OpenStudy (anonymous):

I'm assuming 4 is the base?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\log a + \log b= \log ab\] \[loga - \log b = log a/b\]

OpenStudy (anonymous):

x = 128

OpenStudy (anonymous):

\[\log_4 x/2 = 3\]

OpenStudy (anonymous):